# Mass Modeling and Kinematics of Galaxy Clusters in Modified Gravity

### v1

@@ 3.5 Solution’s existence
Finally, in fig. 2, we show the comparison of the screening functions’ behaviour for different mass models. us comparison amongst the mass models shown for the same value of $\frac{\phi_{\infty}}{\mathcal{B}}$, we can immediately notice that the $g$NFW profile provides larger $r_{\rm c}$ wrt the Isothermal and Burkert mass profiles. This implies that for the same normalization set by $\frac{\phi_{\infty}}{\mathcal{B}}$, using the $g$NFW mass profile most of the cluster is screened. It is also interesting to note that the dependence of the solution to the screening varies significantly in the Isothermal and Burkert cases in comparison to the $g$NFW case. This in turn implies that the $g$NFW profile coupled with the chameleon field screens the fifth force effects within the galaxy cluster more effectively than the two other mass profiles in comparison here. For the purpose of illustration here we show the $g$NFW profile assuming $\gamma=1.2$.
@@ 4 Comparison with numerical solutions
The set of numerical solutions of Eq. 2.4 were obtained through a $6^{th}$-order explicit Runge-Kutta integrator while the appropriate boundary conditions – $\nabla^{2}\phi\approx 0$ at the centre of the mass distribution, and $d\phi/dr=0$ at infinity – were imposed through a Newton-Rapshon shooting method. In order to avoid the divergence at the centre of the mass density distribution associated with some of the models, an inner cutoff radius was imposed. The value of the latter ranged between $5-10\%$ of the screening radius in order to get the best fit for the semi-analytic approach.
Comparative results between the semi-analytic (solid) and the full numerical solutions (dashed) can be seen in Fig. 3 for all the mass density models under consideration, where $r_{s}=0.5$ and $\beta=0.5$.
Figure 3: Semi-analytic approximation for the effective mass $M_{\text{eff}}(r)$ (solid lines) compared with the numerical solution (points) for the different mass ansatz defined in Section 2.2. The bottom plot indicates the relative difference between the two. The parameters adopted are $\rho_{s}=5\times 10^{14}\,\text{M}_{\odot}/\text{Mpc}^{3}$, $r_{s}=0.5\,\text{Mpc}$, $\beta=0.5$, $\phi_{\infty}=5\times 10^{-5}$.
As can be seen, besides some slight differences in the screening radius transition and asymptotic behaviour (where the strongest assumptions and approximations were made), the semi-analytic approximation describes with a high degree of agreement the full numerically obtained results, giving confidence for their use in more complex calculations that will proceed.
@@ 5 Kinematics mass reconstruction: simulations and testing
Using an upgraded version of the ClusterGEN code (see [65]), we generate synthetic, spherically symmetric, isolated systems of collisionless particles in dynamical equilibrium. The particles are distributed according to an NFW number density profile $\nu(r)=\nu_{\text{NFW}}(r,r_{\nu})$, with a characteristic scale radius $r_{\nu}=0.5\,\text{Mpc}$ [^7] The distribution of the member galaxies in clusters can be different from that of the total matter density, which accounts for several mass components (see e.g. [67, 68]). Here, the NFW ansatz for simplicity, but no significant changes have been found when adopting other density models.. The velocity field follows a dispersion along the radial direction, $\sigma^{2}_{r}$, obtained by solving the Jeans’ equation:
@@
where $r_{\beta}$ is a scale radius equivalent to the scale radius $r_{-2}$ of the assumed mass model, and the value of the anisotropy at infinity, $\eta_{\infty}$, is set to $0.5$. eq. 5.2 is found to provide an adequate fit for the velocity anisotropy of cluster-size halos in cosmological simulations (e.g. [70]).
@@
which can be connected to the scale radius of each density profile in eqs. 2.5, 2.7, 2.6, 2.8, 2.9 and 2.10. In particular, $r_{-2}=r_{\text{s}}$ for the NFW and Einasto models, while $r_{-2}\simeq 1.52\,r_{\text{s}}$ for Burkert, $r_{-2}=\sqrt{3}\,r_{\text{s}}$ for Isothermal, $r_{-2}=r_{\text{s}}/(b-1)$, for b-NFW and $r_{-2}=(2-\gamma)r_{\text{s}}$ for $g$NFW. For simplicity, assume $r_{\beta}=r_{-2}$ in the velocity anisotropy profile.
The total gravitational potential of our synthetic clusters is then reconstructed by means of MG-MAMPOSSt [^9] Code is publicly available at https://github.com/Pizzuti92/MG-MAMPOSSt. program package (see [65, 52]), a modified version of the MAMPOSSt code [72], which determines the mass, anisotropy and number density profiles of spherically symmetric systems by solving the Jeans’ analyses. The code performs a maximum likelihood fit of the line-of-sight velocity field, $v_{z}$, and the projected positions, $R$, of the member galaxies, given parametric input models of anisotropy profile ($\eta(r)$), number density ($\nu(r)$), and the gravitational potential ($\Phi(r)$). The dataset $(R,v_{z})$ is called projected phase space (pps), which is our final simulated observable. Figure 4 shows the pps for the cluster generated with the NFW profile in GR (left) and in Chameleon gravity (right). As one can already visually inspect, the velocity dispersion of the cluster in modified gravity scenario tends to be larger than the one in GR, due to the additional contribution of the fifth force.
@@
Let us also consider an unbiased determination of the number density profile $\nu(r)$, which can be generally obtained externally from MG-MAMPOSSt by fitting the observed projected distribution of galaxies in clusters (see e.g. [73, 74]). As for the "lensing" prior on $r_{200}$ and $r_{\text{s}}$, as done in [52], assume a Gaussian distribution centred on the true values $P(r_{200}^{t},r_{\text{s}}^{t})$, with standard deviations based on reliable uncertainties, given current mass lensing estimates: $\sigma_{r_{200}}=0.1\,r_{200}$ and $\sigma_{r_{\rm s}}=0.3\,r_{\text{s}}$. The correlation is set to $\mu=0.5$, checking that varying this value produces irrelevant changes in the results.
@@
The aim of this section is twofold: a) determine whether a cluster, generated in $\Lambda$CDM, produces spurious detection of modified gravity when the mass is reconstructed in a Chameleon gravity framework assuming the wrong model for the matter density, hereafter referred to as case I. And b) investigate how the shape of the total mass distribution influences the fifth force imprint of a cluster-size halo in a Chameleon universe in the joint kinematics+lensing analyses (aka case II). Note that, in both cases, possible effects induced by the halo’s triaxiality are ignored, and solely spherical symmetry is considered. Results are presented in terms of the re-scaled variables $\mathcal{Q}_{2}=\beta/(1+\beta)$ and $\phi_{2}=1-\exp\left[\phi/(10^{4}\,M_{\rm Pl})\right]$, which spawn the range $[0,1]$. fig. 5 shows the marginalized distributions of $\mathcal{Q}_{2}$ and $\phi_{2}$ in case I. The darker and lighter shaded areas indicate the one and two-$\sigma$ regions allowed in the parameter space. No significant shifts towards $\phi_{2}\neq 0$ and $\mathcal{Q}_{2}\neq 0$ are observed when a “wrong” mass model is used to estimate the gravitational potential independently of the mass density model. This indicates a sub-dominance of the possible biases due to the mass modelling when compared to the statistical uncertainties, suggesting the robustness of our method against these specific systematics. Nevertheless, the relation between the efficiency of the screening mechanism and the chosen density profile is clearly visible in the shape of the distributions. Models like Burkert, Isothermal or Einasto tend to have a smaller screening radius for the same value of $r_{200}$, $r_{-2}$ and Chameleon parameters, resulting in tighter constraints in the space $(\phi_{2},\mathcal{Q}_{2})$, due to the additional un-screened region. Note that the same results have been found for clusters generated with other mass distributions besides the NFW.
@@
It is worth noticing the difference between the halo generated assuming a Burkert profile (solid lines and filled regions in fig. 6) and the one generated using an NFW model (dashed lines/contours). As already mentioned above, the effectiveness of the screening strongly depends on the assumed mass model, given the same background modification to gravity.
@@ 5.1 Bayesian evidence
Note, however, that the assessment so far is done for case II when simulating the galaxy clusters in a strongly modified gravity regime. A similar comparison can be performed for the GR clusters simulated in case I. Even in this case, we find that the Bayesian evidence shows no bias for a spurious detection when an incorrect mass model is assumed in assessing the constraints on the Chameleon field. For all the mass models assumed here, we find $\log(\mathcal{B})<0$, implying the NFW model is always preferred and a maximum disadvantage is found for the Burkert model with $\log(\mathcal{B})\sim-1.2$.
Figure 5: Marginalized one and two-dimensional distributions of $\phi_{2}$ and $\mathcal{Q}_{2}$ reconstructed by applying the MG-MAMPOSSt method to the NFW-GR cluster (case I). Each plot refers to a different mass density ansatz used in the fit. Inner and outer shaded regions in the contour plots refer to one and two-$\sigma$ contours. The plots are made by using the getdist package, see also [78].
[…]

### v2

@@ 3.5 Solution’s existence
Finally, in Fig 3, we show the comparison of the screening functions’ behaviour for different mass models. Let us compare amongst the mass models shown for the same value of $\frac{\phi_{\infty}}{\mathcal{B}}$; we can immediately notice that the $g$NFW profile provides larger $r_{\rm c}$ with respect to the Isothermal and Burkert mass profiles. This implies that for the same normalization set by $\frac{\phi_{\infty}}{\mathcal{B}}$, using the $g$NFW mass profile more of the cluster is screened. It is also interesting to note that the dependence of the solution to the screening varies significantly in the Isothermal and Burkert cases in comparison to the $g$NFW case.
This, in turn, implies that the $g$NFW profile coupled with the chameleon field screens the fifth force effects within the galaxy cluster more effectively than the two other mass profiles in comparison here. For the purpose of illustration, here we show the $g$NFW profile assuming $\gamma=1.2$. Note that $r_{c}$ grows with $\gamma$, and it remains larger than that obtained in the case of Burkert or Isothermal profiles for the same value of $\phi_{\infty}/\mathcal{B}$ down to the lower limit of $\gamma=0$. This further indicates that the difference in $r_{c}$ will be even larger for $\gamma\to 2$ in comparison to $\gamma=1.2$.
@@ 4 Comparison with numerical solutions
The set of numerical solutions of Eq. 2.4 were obtained through a $6^{th}$-order explicit Runge-Kutta integrator while the appropriate boundary conditions – $\nabla^{2}\phi\approx 0$ at the centre of the mass distribution, and $\text{d}\phi/\text{d}r=0$ at infinity – were imposed through a Newton-Rapshon shooting method. In order to avoid the divergence at the centre of the mass density distribution associated with some of the models, an inner cutoff radius was imposed. The value of the latter ranged between $5-10\%$ of the screening radius in order to get the best fit for the semi-analytic approach. The appropriate boundary condition at infinity was imposed by considering a numerically small value of the scalar field derivative, $\sim 10^{-8}$, at a scaled radius $x$ several times larger than the main mass distribution, $x_{max}\approx 10^{3}$.
Comparative results between the semi-analytic (solid) and the full numerical solutions (points) can be seen in Fig. 4 for all the mass density models under consideration, assuming $\rho_{s}=5\times 10^{14}\,\text{M}_{\odot}/\text{Mpc}^{3}$, $r_{s}=0.5\,\text{Mpc}\,$, $\beta=0.5$ and $\phi_{\infty}=5\times 10^{-5}$.
Figure 4: Left: Semi-analytic approximation for the effective mass $M_{\text{eff}}$ (solid lines) as a function of $x=r/r_{\rm s}$ compared with the numerical solution (points) for the different mass ansatz defined in Section 2.2. The bottom plot indicates the relative difference between the two. Right: the same for the total dynamical mass $M_{\text{tot}}=M+M_{\text{eff}}$. The adopted parameters are $\rho_{s}=5\times 10^{14}\,\text{M}_{\odot}/\text{Mpc}^{3}$, $r_{s}=0.5\,\text{Mpc}$, $\beta=0.5$, $\phi_{\infty}=5\times 10^{-5}$.
Let us now quickly analyse the difference between the semi-analytical solution and the full non-linear numerically obtained solution, Fig. 4. Both the effective mass (top left) and the total mass (top right), as well as the relative difference between the semi-analytically obtained and the numerically obtained solutions (bottom), are represented in Fig. 4. For the study of how well the semi-analytical solution describes the true solutions, let us analyse the relative difference for the effective mass. As $x$ increases from the origin $x=0$ until the defined cut-off radius, $x=x_{\text{s}}$ – where the strongest assumptions and approximations were made –, while the semi-analytical is set to zero, the numerical solutions are small but non-zero, originating a large relative difference which is accentuated at the transitional scaled radius $x_{\text{s}}$, where the numerical starts to gain significant non-zero values before the semi-analytical (smoother transition from a negligible value). This behaviour is, however, cancelled as one goes away from the mass distribution (increase $x$) to the background configuration (no mass distribution, flat scalar field profile). At this point, the semi-analytical and the numerically obtained solutions coincide almost perfectly, with a maximum relative error of $10^{-3}$. Hence, besides some slight differences in the screening radius transition and asymptotic behaviour, the semi-analytic approximation describes with a high degree of agreement the full numerically obtained results, giving confidence for their use in more complex calculations that will proceed. Note that the maximum discrepancy between the numerical and semi-analytic approach in the total mass $M(r)+M_{\text{eff}}(r)$ (right plots) is of $7\%$, and corresponds to $x=\sim 2$ (i.e. $r\sim 1\,\text{Mpc}$ for the adopted value of $r_{\rm s}$).
@@ 5 Kinematics mass reconstruction: simulations and testing
Using an upgraded version of the ClusterGEN code (see [66]), we generate synthetic, spherically symmetric, isolated systems of collisionless particles in dynamical equilibrium. The particles are distributed according to an NFW number density profile $\nu(r)=\nu_{\text{NFW}}(r,r_{\nu})$, with a characteristic scale radius $r_{\nu}=0.5\,\text{Mpc}$ [^7] The distribution of the member galaxies in clusters can be different from that of the total matter density, which accounts for several mass components (see e.g. [68, 69]). Here, we consider the NFW ansatz for simplicity, but no significant changes have been found when adopting other density models.. The velocity field follows a dispersion along the radial direction, $\sigma^{2}_{r}$, obtained by solving the Jeans’ equation:
@@
where $r_{\beta}$ is a scale radius equivalent to $r_{-2}$ of the assumed mass model, and the value of the anisotropy at infinity, $\eta_{\infty}$, is set to $0.5$. Eq. 5.2 is found to provide an adequate fit for the velocity anisotropy of cluster-size halos in cosmological simulations (e.g. [71]).
@@
which can be connected to the scale radius, $r_{-2}$, of each density profile in Eqs. 2.5, 2.7, 2.6, 2.8, 2.9 and 2.10. In particular, $r_{-2}=r_{\text{s}}$ for the NFW and Einasto models, while $r_{-2}\simeq 1.52\,r_{\text{s}}$ for Burkert, $r_{-2}=\sqrt{3}\,r_{\text{s}}$ for Isothermal, $r_{-2}=r_{\text{s}}/(b-1)$, for b-NFW and $r_{-2}=(2-\gamma)r_{\text{s}}$ for $g$NFW.
The total gravitational potential of our synthetic clusters is then reconstructed by means of MG-MAMPOSSt [^9] Code is publicly available at https://github.com/Pizzuti92/MG-MAMPOSSt. program package (see [66, 53]), a modified version of the MAMPOSSt code [73], which determines the mass, anisotropy and number density profiles of spherically symmetric systems by solving the Jeans’ analyses. Given parametric input models of anisotropy profile ($\eta(r)$), number density ($\nu(r)$), and the gravitational potential ($\Phi(r)$), the code performs a Monte Carlo Markov-Chain (MCMC) sampling of the parameter space, using as input data the line-of-sight velocity field, $v_{z}$, and the projected positions, $R$, of the member galaxies. The dataset $(R,v_{z})$ is called projected phase space (pps), which is our final simulated observable. In particular, MG-MAMPOSSt computes the probability $q(R_{i},v_{z,i}|\theta)$ of finding a galaxy at the point $R_{i},v_{z,i}$ in the pps. The (log) likelihood is then given by $\mathcal{L}_{\text{dyn}}(\theta)=\sum_{i}\text{ln}q(R_{i},v_{z,i}|\theta)$.
Figure 5 shows the pps for the cluster generated with the NFW profile in GR (left) and in Chameleon gravity (right). As one can already visually inspect, the velocity dispersion of the cluster in modified gravity scenario tends to be larger than the one in GR, due to the additional contribution of the fifth force.
@@
MG-MAMPOSSt requires also information about the number density profile $\nu(r)$ of the galaxies, which can be generally obtained externally from MG-MAMPOSSt by fitting the observed projected distribution of galaxies in clusters (see e.g. [74, 75]). Thus, the scale radius $r_{\nu}$ of the number density profile can be considered as a fixed parameter in the MCMC, with the value provided by the external fit; note that this is independent of the model adopted for the total mass profile in the Jeans’ equation. Here by construction, the mock pps are built in such a way fitting the number density of galaxies with a NFW model will provide the correct value of the scale radius $r_{\nu}=0.5\,\text{Mpc}$. It is important to mention that variation of the constraints on the inferred dynamical mass in modified gravity, induced by variation of the number density scale radius $r_{\nu}$ are negligible (see e.g. [48]).
As shown in [66], kinematics data alone cannot constrain efficiently the parameter space of Chameleon models, due to the degeneracy between the parameters of the density profile $r_{200}$ and $r_{\text{s}}$ and the fifth force. As such, additional independent information should be provided on $r_{200}$ and $r_{\text{s}}$ in order to break the degeneracy. In this regard, cluster mass determination with gravitational lensing analyses is a powerful tool to be combined with kinematic data. Indeed, due to the conformal structure of Chameleon theories and conformal invariance of null geodesics, photons do not perceive the effect of the fifth force (see e.g. [27]). We thus simulate the availability of a lensing-like information as a Gaussian distribution $P_{\text{l}}=\mathcal{G}(r_{200},r_{\text{s}})$. Despite all effects due to asphericity are neglected, as well as systematics in the lensing modeling (see [76] for a review), to determine the mean values $\bar{r}_{200},\bar{r}_{\text{s}}$ we should still account for the effect of choosing the wrong mass profile also in a lensing analysis. Indeed, even if the value of $r_{200}$ should be, in principle, the same when inferred by the wrong mass model (as this does not depend on the shape of the halo), the scale radius $r_{\text{s}}$ differs from profile to profile. Given the value of $r_{-2}$, which is fixed by the "true" mass profile, each mass model would infer a $r_{\text{s}}$ which is, in general, a function of $r_{-2}$ as discussed above. Thus, the central value of $r_{\text{s}}$ in the lensing Gaussian varies for each profile according to $\bar{r}_{\text{s}}=f(r_{-2}=0.5\,\text{Mpc})$.
As for the standard deviations, those are based on reliable uncertainties given current mass lensing estimates ([77]): $\sigma_{r_{200}}=0.1\,r_{200}$ and $\sigma_{r_{\rm s}}=0.3\,r_{\text{s}}$. One can also check that variations of the correlation do not produce significant changes in the results. With this into account, let us set the correlation as $\mu=0.5$.
@@
The aim of this section is twofold: a) determine whether a cluster, generated in $\Lambda$CDM, produces spurious detection of modified gravity when the mass is reconstructed in a Chameleon gravity framework assuming the wrong model for the matter density, hereafter referred to as case I. And b) investigate how the shape of the total mass distribution influences the fifth force imprint of a cluster-size halo in a Chameleon universe in the joint kinematics+lensing analyses (aka case II). Note that, in both cases, possible effects induced by the halo’s triaxiality are ignored, and solely spherical symmetry is considered. Results are presented in terms of the re-scaled variables $\mathcal{Q}_{2}=\beta/(1+\beta)$ and $\phi_{2}=1-\exp\left[\phi/(10^{4}\,M_{\rm Pl})\right]$, which spawn the range $[0,1]$. Within this range, we consider uninformative uniform prior for both parameters, as already done in previous works (see e.g. [46, 51]). As for the anisotropy parameter $\eta_{\infty}$, as a standard procedure (e.g. [74]) we consider flat priors in the range $[0.5,7]$ in the parameter $\mathcal{A}_{\infty}=(1-\eta_{\infty})^{-1/2}$.
A sampling of 110000 points of the joint log-likelihood
$\mathcal{L}_{\text{tot}}=\mathcal{L}_{\text{dyn}}+\ln[P(r_{200},r_{\text{s}})]\,,$ (5.4)
in the parameter space, assuming the above mentioned priors, is performed through a Metropolis-Hastings algorithm. We discard the first 10000 points as a burn-in phase, resulting in a final chain of 100000 samples. The convergence is ensured for each run by performing $n=5$ chains and computing the corresponding Gelman-Rubin diagnostic coefficients $\hat{R}$ [78], checking that the requirement $\hat{R}\lesssim 1.1$ is always satisfied.
Fig. 6 shows the marginalized distributions of $\mathcal{Q}_{2}$ and $\phi_{2}$ in case I. The darker and lighter shaded areas indicate the one and two-$\sigma$ regions allowed in the parameter space for the Burkert-generated halo while the inner and outer dashed contours are the same for the NFW-generated halo. No significant shifts towards $\phi_{2}\neq 0$ and $\mathcal{Q}_{2}\neq 0$ are observed when a “wrong” mass model is used to estimate the gravitational potential independently of the mass density model. This indicates a sub-dominance of the possible biases due to the mass modelling when compared to the statistical uncertainties, suggesting the robustness of our method against these specific systematics. Nevertheless, the relation between the efficiency of the screening mechanism and the chosen density profile is clearly visible in the shape of the distributions. Models like Burkert, Isothermal or Einasto tend to have a smaller screening radius for the same value of $r_{200}$, $r_{-2}$ and Chameleon parameters, resulting in tighter constraints in the space $(\phi_{2},\mathcal{Q}_{2})$, due to the additional un-screened region. Note that the same results have been found for clusters generated with other mass distributions besides the NFW and Burkert models.
Figure 6: Marginalized one and two-dimensional distributions of $\phi_{2}$ and $\mathcal{Q}_{2}$ reconstructed by applying the MG-MAMPOSSt method to the GR clusters (case I). Shaded regions refer to a cluster generated with a Burkert profile, while the dashed contours are for a cluster generated assuming a NFW model. Each plot refers to a different mass density ansatz used in the MG-MAMPOSSt fit. Inner and outer shaded regions (or inner and outer dashed contour lines) in the contour plots refer to one and 1-$\sigma$ contours. The plots are made by using the getdist package, see also [79].
Figure 7: Marginalized posteriors for the one and two-dimensional distributions of $\phi_{2}$ and $\mathcal{Q}_{2}$ from the MG-MAMPOSSt analysis of two clusters generated in Chameleon gravity (case II) assuming a Burkert profile (solid shaded areas) and a NFW profile (dashed contours). Each plot refers to a different mass density ansatz used in the fit. Inner and outer shaded regions, as well as inner and other dashed contours, refer to 1-$\sigma$ and 2-$\sigma$ limits, respectively. The corresponding parameter constraints are presented in table 1. The simulated cluster is represented by * in each panel, corresponding to $\{\phi_{2},\mathcal{Q}_{2}\}=\{0.63,0.5\}$.
@@
In both cases, it is worth noticing the difference in the marginalized contours between the halos generated assuming a Burkert profile (solid lines and filled regions in Figs. 6, 7) and the one generated using a NFW model (dashed lines/contours). In particular, the NFW model generally provides wider constraints on the Chameleon parameters; as already mentioned above, this comes as a consequence of the effectiveness of the screening, which strongly depends on the assumed mass model, given the same background universe (GR or modified gravity).
In Fig. 8 a list of plots of the reconstructed field profiles for case II is further shown. As expected, the stronger screening efficiency of the NFW model results in wider uncertainties in the field profile. For large radii, all profiles tend to the true background value $\phi/M_{\rm Pl}=10^{-4}c^{2}$, with some variability. It is worth mentioning that the best agreement is found when the true profile is used in the MG-MAMPOSSt analysis.
Figure 8: Radial field profiles $\phi/M_{\rm Pl}$, in unit of $c^{2}$, computed from the MG-MAMPOSSt posteriors of case II. Solid lines: Burkert-generated halo. Dashed line NFW-generated halo. The shaded areas (dashed top and lower lines) represent the $1\,\sigma$ regions.
@@ 5.1 Bayesian evidence
Note, however, that the assessment so far is done for case II when simulating the galaxy clusters in a strongly modified gravity regime. A similar comparison can be performed for the GR clusters simulated in case I. Even in this case, we find that the Bayesian evidence shows no bias for a spurious detection when an incorrect mass model is assumed in assessing the constraints on the Chameleon field. For the NFW-generated halo all the mass models assumed have $\log(\mathcal{B})<0$, implying the NFW model is always preferred and a maximum disadvantage is found for the Burkert model with $\log(\mathcal{B})\sim-1.2$. When the halo is generated with the Burkert model, a slight preference ($\log(\mathcal{B})\sim 1.0$) is found for the Hernquist fitting model, but this is statistically not relevant.